There are 3 less than twice as many cats than dogs in the local shelter. If there are 210 cats and dogs combined, how many dogs are in the animal shelter?

Answer by checkley79(3341)

D=2C-3
D+C=210
(2C-3)+C=210
2C+C=210+3
3C=213
C=213/3
C=71 NUMBER OF CATS.
D=2*71-3
D=142-3=139 NUMBER OF DOGS.
PROOF:
139+71=210
210=210